OpenStudy (anonymous):

what is the volume of a regular square pyramid with the base sides of 10 cm and a height of 12cm

jimthompson5910 (jim_thompson5910):

V = (s^2*h)/3 V = (10^2*12)/3 V = (100*12)/3 I'll let you finish

OpenStudy (lilai3):

congrats on your first question, @jaetiara!